Work in Process Account Data for Two Months; Cost of Production Reports

Pittsburgh Aluminum Company uses a process cost system to record the costs of manufacturing rolled aluminum, which consists of the smelting and rolling processes. Materials are entered from smelting at the beginning of the rolling process. The inventory of Work in Process—Rolling on September 1 and debits to the account during September were as follows:

Bal., 800 units, 80% completed:
Direct materials (800 x $4.6) $ 3,680
Conversion (800 x 80% x $1.9) 1,216
$ 4,896
From Smelting Department, 18,480 units $86,856
Direct labor 22,503
Factory overhead 12,117

During September, 800 units in process on September 1 were completed, and of the 18,480 units entering the department, all were completed except 1,900 units that were 30% completed. Charges to Work in Process—Rolling for October were as follows:

From Smelting Department, 21,300 units $104,370
Direct labor 30,070
Factory overhead 16,193

During October, the units in process at the beginning of the month were completed, and of the 21,300 units entering the department, all were completed except 1,000 units that were 40% completed.

Required:

1. Enter the balance as of September 1 in a four-column account for Work in Process—Rolling. Record the debits and the credits in the account for September. Construct a cost of production report and present computations for determining (a) equivalent units of production for materials and conversion, (b) costs per equivalent unit, (c) cost of goods finished, differentiating between units started in the prior period and units started and finished in September, and (d) work in process inventory. If an amount box does not require an entry, leave it blank.

2. Provide the same information for October by recording the October transactions in the four-column work in process account. Construct a cost of production report, and present the October computations (a through d) listed in part (1). If an amount box does not require an entry, leave it blank.

ACCOUNT Work in Process-Rolling Department ACCOUNT NO.
Balance
DATE ITEM POST. REF. DEBIT CREDIT DEBIT CREDIT
October 1 Balance
October 31 Smelting Dept., 21,300 units at $4.9
October 31 Direct labor
October 31 Factory overhead
October 31 Finished goods
October 31 Bal., 1,000 units, 40% completed

If an amount is zero, enter in a zero "0". Round cost per unit answers to the nearest cent.

Pittsburgh Aluminum Company
Cost of Production Report-Rolling Department
For the Month Ended October 31
Whole Units Equivalent Units
Units Direct Materials (a) Conversion (a)
Units charged to production:
Inventory in process, October 1
Received from Smelting Department
Total units accounted for by the Rolling Department
Units to be assigned costs:
Inventory in process, October 1
Started and completed in October
Transferred to finished goods in October
Inventory in process, October 31
Total units to be assigned costs
Costs
Costs Direct Materials Conversion Total Costs
Cost per equivalent unit:
Total costs for October in Rolling Department $ $
Total equivalent units
Cost per equivalent unit (b) $ $
Costs assigned to production:
Inventory in process, October 1 $
Costs incurred in October
Total costs accounted for by the Rolling Department $
Costs allocated to completed and partially completed units:
Inventory in process, October 1 balance (c) $
To complete inventory in process, October 1 (c) $ $
Cost of completed October 1 work in process $
Started and completed in October (c)
Transferred to finished goods in October (c) $
Inventory in process, October 31 (d)
Total costs assigned by the Rolling Department $

3. The cost per equivalent unit for direct materials increased  from August to October. The cost per equivalent unit for conversion costs increased  from August to October. These changes should  be investigated for their underlying causes, and any necessary corrective actions should be taken.
Answer #1

Part 1

PITTSBURGH ALUMINUM COMPANY

Cost of Production Report—Rolling Department

For the Month Ended September 30

UNITS

Equivalent Units

Whole Unit

Direct Materials

Conversion

Units charged to production:

Inventory in process, September 1

800

Received from Smelting Department

18480

Total units accounted for by the Rolling Department

19280

Units to be assigned cost:

Inventory in process, September 1 (80% completed)

800

0

160

Started and completed in September

16580

16580

16580

Transferred to finished goods in September

17380

16580

16740

Inventory in process, September 30 (30% complete)

1900

1900

570

Total units to be assigned cost

19280

18480

17310

COSTS

Costs

Direct Materials

Conversion

Total

Costs per equivalent unit:

Total costs for September in Rolling Department

86856

34620

Total equivalent units

18480

17310

Cost per equivalent unit

4.70

2.00

Costs assigned to production:

Inventory in process, September 1

4896

Costs incurred in September

121476

Total costs accounted for by the Rolling Department

126372

Cost allocated to completed and partially completed units:

Inventory in process, September 1 balance

4896

To complete inventory in process, September 1

0

320

320

Cost of completed September 1 work in process

5216

Started and completed in September

77926

33160

111086

Transferred to finished goods in September

116302

Inventory in process, September 30

8930

1140

10070

Total costs assigned by the Rolling Department

126372

22503+12117 =34620

160*2= 320

16580*4.70 = 77926

16580*2 = 33160

1900*4.70 = 8930

570*2.00= 2888

PITTSBURGH ALUMINUM COMPANY

Cost of Production Report—Rolling Department

For the Month Ended October 31

UNITS

Equivalent Units

Whole Unit

Direct Materials

Conversion

Units charged to production:

Inventory in process, October 1

1900

Received from Smelting Department

21300

Total units accounted for by the Rolling Department

23200

Units to be assigned cost:

Inventory in process, October 1 (30% completed)

1900

0

1330

Started and completed in September

20300

20300

20300

Transferred to finished goods in September

22200

20300

21630

Inventory in process, October 30 (40% complete)

1000

1000

400

Total units to be assigned cost

23200

21300

22030

COSTS

Costs

Direct Materials

Conversion

Total

Costs per equivalent unit:

Total costs for October in Rolling Department

104370

46263

Total equivalent units

21300

22030

Cost per equivalent unit

4.90

2.10

Costs assigned to production:

Inventory in process, October 1

10070

Costs incurred in October

150633

Total costs accounted for by the Rolling Department

160703

Cost allocated to completed and partially completed units:

Inventory in process, October 1 balance

10070

To complete inventory in process, October 1

0

320

2793

Cost of completed October 1 work in process

12863

Started and completed in September

99470

42630

142100

Transferred to finished goods in September

154963

Inventory in process, September 30

4900

840

5740

Total costs assigned by the Rolling Department

160703

30070+16193 = 46263

1330*2.10 = 2793

20300*4.90 = 99470

20300*2.10 = 42630

1000*4.90 = 4900

400*2.10 = 840

Part 3

The cost per equivalent unit for direct materials increased from $4.60 in August to $4.70 in September to $4.90 in October. The cost per equivalent unit for conversion costs increased from 1.90 in August to $2.00 in September to $2.10 in October. These increases should be investigated for their underlying causes, and any necessary corrective actions should be taken.